# The Illusive Slump of Disruptive Patents

### v1

@@ 4 Assessing Truncation Effects
Fig. 3 categorizes focal patents by the number of backward citations to demonstrate the impact of truncation. The left panel uses the Park et al. (2023) methodology; the right panel uses our methodology.
Figure 3: Patent Percentage by Backward Citation Categories
Notes: The figure divides patents into five categories based on the number of backward citations. The vertical-axis shows the percentage of focal patents in each category. The left panel (Park et al. (2023) Methodology) excludes backward citations to patent grants published before 1976; the right panel (No Truncation Methodology) includes backward citations to patent grants regardless of publication year.
Both panels show that the majority of patents have backward citations. However, notable differences are found depending upon whether truncation is or is not present. The left (truncation) panel indicates that 38 percent of patents had no backward citations in 1980 and 99 percent of patents had no backward citations in 1976 (the truncation year), given that the only backward citations included are those patents also published in 1976. The right (no truncation) panel indicates that less than two percent of patents had no backward citations in these years. Fig. 3 also indicates that patents published further from the truncation year have lower proportions of truncated backward citations. The reason is again straightforward: More recent patents tend to cite more recent patents.
Figure 4: Truncation Time is Not Important
@@
It is readily apparent that the number of years considered for backward citations matters. The analysis of Park et al. (2023) begins in 1980 and includes backward citations to 1976 (the first year that USPTO patents are available). Hence, patents published in 1980 include four years of backward citations. Fig. 4 shows that had their analysis started in 1976 (i.e., the year of truncation), the bias would be much larger: nearly all patents from 1976 would be classified as disruptive. Fig. 4 also shows that the starting year decision does not alleviate truncation bias and measurement concerns: e.g., had their analysis started in 1990 or in 2000, a four-year window for backward citations would produce similarly sharp declines in the CD index in the earliest years after the truncation year which then stabilizes and converges to the unbiased index over time.
Fig. 5 indicates that this bias is not identical across the technologies examined by Park et al. (2023). The bias is most pronounced in the aggregate technology "Mechanical" and least pronounced in "IT". The difference is again most likely explained by truncation: IT patents have developed more recently and rapidly in comparison to Mechanical patents, suggesting older patents are relatively more often cited in the latter category than in the former category (see Appendix Fig. B.4). As a result, fewer backward citations are truncated in IT patents and more backward citations are truncated in Mechanical patents.
Figure 5: Size of Truncation Bias Differs by Technologies
The figure shows the average $CD_{5}$ index for various aggregate WIPO technologies, as determined in Table B.3. The blue line shows the CD index when backward citations to granted U.S. patents published before 1976 are excluded, per the Park et al. (2023) methodology. The yellow line shows the $CD_{5}$ index when all backward citations are included, per our methodology.
@@ 5 Updating Years and Patent Law Change
We examine two factors that could affect the CD index post-2010. First, we update the analysis window using patent data to 2016. Second, we consider a major change to U.S. patent law: in particular, the Inventor Protection Act of 1999 requires that U.S. patent applications be published 18 months after the initial application filing, effectively shifting citation patterns from strictly patent grants to both patent grants and patent applications Johnson and Popp (2003). [^4] See Appendix A for more details on the additional data used, and Appendix B for a more detailed discussion of how excluding citations to patent applications affects the CD index of patents. Fig. 6 shows these post-2010 changes in three average $CD_{5}$ indices: the Park et al. (2023) methodology (I); our methodology including all backward citations to strictly patent grants (II); and our methodology including all backward citations to patent grants and patent applications (III). Methodology (III) considers only those citations to patent applications that receive patent grants by the end of 2021. This approach best achieves consistency between the pre- and post-Inventor Protection Act periods, as only patent grants were allowable citations. [^5] Fig. B.5 in the Appendix shows similar results when citations to patent applications that have not (yet) received patent grants are included.
@@
Figure 6: Patent Law Change Creates Exclusion Bias in Park et al. (2023) Methodology
@@
The figure shows the average $CD_{5}$ index under different methodologies: the Park et al. (2023) methodology that excludes backward citations to patent grants before 1976 (I); our methodology that includes all backward citations to strictly patent grants (II); and our methodology that includes all backward citations to patent grants and patent applications (III).
@@
Fig. 6 is identical to Fig. 1 up to 2001 because—prior to the Inventor Protection Act’s passage—patents only cited patent grants. The inclusion of citations before 1976 eliminates truncation bias and subsequently decreases the CD index over 1980-2005—as seen by comparing (I) to (II) and (III). The inclusion of citations to patent applications eliminates exclusion bias and subsequently decreases the CD index over 2005-2016—as seen by comparing (I) and (II) to (III). Correcting for both biases indicates that the average CD index decreases slightly over 1980-2016 and is relatively stable since 2005—as shown in methodology (III).
@@ 6 Increasing Number of Highly Disruptive Patents
[…]
Figure 7: Investigation of Highly Disruptive Patents
(a) Different Calculation Methods
(b) Method (III) and Different Technologies
The figure illustrates the number of and change in highly disruptive patents (i.e., $CD_{5}\in(0.75,1]$) over time. Panel (a) shows the number of highly disruptive patents under different methodologies: the Park et al. (2023) methodology that excludes backward citations to patent grants before 1976 (I); our methodology that includes all backward citations to strictly patent grants (II); and our methodology that includes all backward citations to patent grants and patent applications (III). The total number of highly disruptive patents differs from that reported in Fig. 4b (p. 141) of Park et al. (2023) because they only consider a subset of technologies. (See Appendix A for more detailed discussion.) Panel (b) shows the evolution of highly-disruptive patents across aggregated WIPO technology categories using methodology (III), normalized using 1980 as a baseline.
Using 1980 as a benchmark, Panel (b) of Fig. 7(b) shows that disruption is heterogeneous across technologies: the number of highly-disruptive IT patents increases nearly twenty-fold; the number of highly-disruptive Pharma patents increases more than four-fold; and the number of highly-disruptive Mechanical patents is stable. These results challenge the Park et al. (2023) argument that "despite large increases in scientific productivity, the number of papers and patents with $CD_{5}$ values in the far right tail of the distribution remains nearly constant over time" (p. 140). Our findings are, however, consistent with a recent scientific publication that indicates "a paper published in Nature Park et al. (2023) showed that research is becoming less disruptive. This does not seem to be the case for the next generation of cancer chemotherapy based on the methionine dependence of cancer" (p. 277, Kubota et al., 2023).
@@
Thus far, we have taken an aggregated methodological approach, but this masks two disparities inherent at the individual patent level. First, the truncation of backward citations misclassifies a large proportion of patents as highly disruptive. Appendix Table B.1 illustrates how $CD_{5}$ categories for patents published in 1980 calculated using the Park et al. (2023) (truncation) methodology are distributed across $CD_{5}$ categories using our (no truncation) methodology. As an example, 25 percent of the USPTO patents in 1980 that are considered highly disruptive under the truncation methodology are actually consolidating under the no truncation methodology.
@@
Second, the exclusion of backward citations to patent applications misclassifies a large proportion of recent patents as highly disruptive. Appendix Table B.2 shows how patents in 2016 that belong to a particular $CD_{5}$ category using the Park et al. (2023) (truncation) methodology are distributed across $CD_{5}$ categories using our (no truncation) methodology and accounting for backward citations to patent applications. As an example, 19 percent of the USPTO patents in 2016 that are considered highly disruptive using the Park et al. (2023) methodology are actually consolidating when accounting for both truncation and backward citations to patent applications. Determining the disruptive potential of individual patents thus requires careful consideration to ensure against bias and misclassification. Such bias and misclassification is the case when the data provided by Park et al. (2023) or Funk and Owen-Smith (2017) are used for analysis and decision-making.
@@ 7 Discussion
Our examination of the Park et al. (2023) CD index methodology was prompted by the stark decrease in the share of disruptive innovation in science and technology the authors find in patents over 1980-2010 and in scientific publications over 1945-2010. At least for patents, our analysis instead suggests that the decrease in the CD index is mainly due to the truncation of backward citations, while a modest increase in the CD index post-2005 is mainly driven by the exclusion of backward citations to patent applications brought on by a change in patent law.
Variations of the index proposed in Park et al. (2023) and based upon other research Leydesdorff et al. (2021); Bornmann et al. (2020a) are also affected by truncation bias and exclusion bias when backward citations to granted patents and to patent applications are not appropriately accounted for. Our analysis is important for the research community that may overlook these fundamental issues and work with biased data – either via the CD index or some other measure.
Our analysis highlights, in particular, the dangers of truncation bias from backward citations and exclusion bias from citations to patent applications. Our correction and methodology shows—in comparison to Park et al. (2023)—that: (1) the CD index starting level is significantly lower, modestly declines up to 2005, and is mostly stable since; (2) the starting truncation "date" does not matter—any backward citation "cutoff" in any chosen year can produce similar CD index patterns; (3) truncation bias differences are heterogeneous across technology classes; and (4) the number of highly disruptive patents has increased—particularly since 2010 and in IT.
@@
Our analysis suggests, in general, caution against overt government policy interventions, research policy changes, or substantive industry and firm reorganizations that seek to improve or disrupt the innovation status-quo. If it was accurate that breakthroughs in science and technology have substantially decreased over time, the current organization of the entire science and technology innovation process could be called into question and should be reevaluated. In particular, reforms that seek to improve the creation of disruptive innovation or restart innovation within R&D centers and research labs—e.g., in government agencies, universities and firms—would become paramount.
But our results suggest—at least in the patent data—that the so-called recent slump in disruptive innovation is illusory. An unbiased CD index is instead markedly stable over the past several decades, and highly disruptive patents have increased overall–especially in IT and Pharma. Thus, an interesting question for future research could investigate why these industries have seen increases in disruptive innovation, relative to others. Given our findings, we should finally emphasize that premature changes in the current innovation processes of science and technology might bring with them unintended consequences.
@@ A Data
The study uses data from the February 21, 2023 version of PatentsView, which includes all U.S. patents published between 1976 and September 29, 2022. As in previous research Park et al. (2023); Funk and Owen-Smith (2017), we focus only on USPTO patents. We include PatentsView data that assigns each USPTO patent to technology categories as defined by the World Intellectual Property Organization (WIPO) Schmoch (2008). The specific WIPO technologies used to create our aggregated technology groups are provided in the Table B.3. It is important to note that our technology classification differs from that of Park et al. (2023), who categorize patents based on NBER technology fields that are no longer available.
@@
In Sections 2-4, we use all granted utility patents over 1980-2010. Park et al. (2023) examine granted utility patents over this timeframe and in the NBER technology fields of "Chemical," "Computers and Communications," "Drugs and Medical," "Electrical and Electronic," and "Mechanical." Their data includes 3,046,672 patents and 29,777,375 backward citations (see replication data file $patentsview\_analytical\_df.csv$ of Park et al. (2023)). We examine granted utility patents over the same timeframe, but cover all technology fields. Our data thus includes 3,662,051 patents and 42,617,016 backward citations in the case of no truncation (i.e., as they appear in the original patent documents) and 35,192,186 backward citations in the case of truncation.
@@
In Sections 5-6, we expand the dataset over 1980-2016, which results in 5,324,224 granted utility patents and 74,943,850 backward citations. With the passage of the Inventor Protection Act of 1999, patents beginning in November 2000 include backward citations to patent applications and to patent grants. Thus, in our methodology (III), we include all backward citations to patent grants and to patent applications that were eventually granted using supplemental PatentsView data. We replace all citations to patent applications with the corresponding patent grant information: i.e., the grant number and publication date. This replacement can result in some backward citations with publication dates later than the citing patent, but the backward citation publication date is not relevant to the $CD_{5}$ index methodology. What matters are publication dates of forward citations. We do not include patent applications published within five years after the publication year of a focal patent but with corresponding patent grants published later in the analysis. Finally, we do not include citations to patent applications that have not been granted by the end of 2021 (the latest date available) in the analysis. This approach results in 87,880,452 backward citations.

### v2

@@ 4 Assessing Truncation Effects
It is readily apparent that the number of years considered for backward citations matters. The closer to (further from) the truncation year, the greater (lesser) the share of patents with no backward citations and, thus, the larger (smaller) the CD index. As shown in Fig. 1, had the Park et al. (2023) methodology started in 1976 (i.e., the truncation year), the bias would be larger: viz., nearly all patents from 1976 would be classified as disruptive. Fig. 3 verifies this relationship using different truncation times and demonstrates that the starting year decision for the CD index does not alleviate truncation bias concerns: had the analysis started, e.g., in 1990 (green line) or in 2000 (black line), a four-year truncation window for backward citations would produce similarly sharp CD index declines in the earliest post-truncation years which then stabilize and converge to the unbiased index over time.
Abbildung 3: Truncation Time is Not Important
@@
Fig. 4 indicates that this bias is not identical across the technologies examined by Park et al. (2023): viz., it is most pronounced in "Mechanicaländ least pronounced in ÏT". The difference is again most likely explained by truncation: IT patents have developed more recently and rapidly in comparison to Mechanical patents, suggesting older patents are relatively more often cited in the latter category than in the former category (see Appendix Fig. B.4). As a result, less (more) backward citations are truncated in IT (Mechanical) patents. For comparison, the yellow line represents the $CD_{5}$ index when all backward citations are included (no truncation) and implies relatively small changes over time and relatively minor differences between technologies.
Abbildung 4: Size of Truncation Bias Differs by Technologies
The figure shows average $CD_{5}$ indices for various aggregate WIPO technologies, as determined in Appendix Table B.3. The blue line shows the CD index when backward citations to granted U.S. patents published before 1976 are excluded, per the Park et al. (2023) methodology. The yellow line shows the $CD_{5}$ index when all backward citations are included, per our methodology.
@@ 5 Updating Years and Patent Law Change
We examine two factors that could affect the post-2010 CD index. First, we update the patent data analysis window to 2021 to extend the $CD_{5}$ index to 2016. Second, we consider a major U.S. patent law change: in particular, the Inventor Protection Act of 1999 requires that U.S. patent applications be published 18 months after the initial application filing—effectively shifting citation patterns from strictly patent grants to patent grants and patent applications Johnson and Popp (2003). [^10] See Appendix A for more detailed discussion on the additional data used, and Appendix B for more detailed discussion of how excluding citations to patent applications affects the CD index.
Fig. 5 shows these post-2010 changes in three average $CD_{5}$ indices using the same data: (I) the Park et al. (2023) methodology; (II) our methodology including all backward citations to patent grants only; and (III) our methodology including all backward citations to patent grants and patent applications. Note that methodology (III) considers only those citations to patent applications that receive patent grants within five years after the publication year of a focal patent, as this approach best achieves consistency between the pre- and post-Inventor Protection Act periods. [^11] Appendix Fig. B.5 shows similar results when citations to patent applications that have not (yet) received patent grants are included.
@@
Abbildung 5: Patent Law Change Creates Exclusion Bias in Park et al. (2023) Methodology
@@
The figure shows the average $CD_{5}$ index under different methodologies: (I) the Park et al. (2023) methodology that excludes backward citations to patent grants before 1976; (II) our methodology that includes all backward citations to strictly patent grants; and (III) our methodology that includes (II) and all backward citations to patent grants and patent applications. The 2010 vertical line denotes the end of the Park et al. (2023) analysis.
@@
Fig. 5 is identical to Fig. 1 up to 2001 because—prior to the Inventor Protection Act’s passage—patents only cited patent grants. The inclusion of citations before 1976 eliminates the truncation bias and subsequently decreases the CD index over 1980-2005—as seen by comparing methodology (I) to methodologies (II) and (III). The inclusion of citations to patent applications eliminates the exclusion bias and subsequently decreases the CD index over 2005-2016—as seen by comparing methodologies (I) and (II) to methodology (III). Correcting for both biases, methodology (III) indicates the average CD index declined slightly from the 54.5 percentile in 1980 to the 52 percentile in 2005, and has been relatively stable since. [^12] Appendix Fig. B.6 shows the correlation of $CD_{5}$ values of focal patents, backward citations of focal patents, citations to focal patents, and citations to focal patents and/or their predecessors for the three methodologies over time. The correlation of the $CD_{5}$ values between the different methods is below 0.5 at the beginning (1980) and end (2016) of our analysis: In 1980, this is mainly due to low correlations for the backward citations of the focal patents; In 2016, this is mainly due to low correlations for citations to the focal patents.
Our analyses clearly document that the Park et al. (2023) CD index patent results are significantly influenced by the time window chosen and the biases present. Extending the patent analysis to 2016, moreover, reveals increasing average CD values—especially since 2005—due to the exclusion of citations to patent applications. In addition, methodology (II) in Fig. 5 shows the CD index evolution if backward citations to pre-1976 patents are considered but backward citations to patent applications are not: a pronounced U-shape results with nearly equal CD index averages in 1980 (start point of the green line) and in 2016 (end point of the yellow line).
@@ 6 Increasing Number of Highly Disruptive Patents
Park et al. (2023) emphasize in their analysis that "despite large increases in scientific productivity, the number of papers and patents with $CD_{5}$ values in the far right tail of the distribution remains nearly constant over time"(p. 140). We evaluate this conclusion with respect to patents—after correcting for truncation bias and exclusion bias—by calculating the number of patents in the three highly disruptive patents groups defined by Park et al. (2023): $CD_{5}\in(0.25,\penalty 10000\ 0.5]$, $CD_{5}\in(0.5,\penalty 10000\ 0.75]$, and $CD_{5}\in(0.75,\penalty 10000\ 1]$. Fig. 6 illustrates the results for each patent group in separate panels. In contrast to Park et al. (2023), our analysis reveals a notable increase in the number of highly disruptive patents over time. [^13] Appendix Figure B.7 shows the number of very highly disruptive patents for five aggregated WIPO technologies. Using the Park et al. (2023) methodology, the absolute number of very highly disruptive patents over 1980-2010 has notably decreased in Chemical and Mechanical; moderately increased in Pharmaceuticals; and notably increased in Electronics and IT. Our Methodology III instead indicates increases in the absolute number of very highly disruptive patents over 1980-2010 in all WIPO technologies except Mechanical.
Abbildung 6: Investigation of Highly Disruptive Patents
The figure shows the number of highly disruptive patents over time for three highly disruptive patent groups: Panel (a) uses $CD_{5}\in(0.25,\penalty 10000\ 0.5]$; Panel (b) uses $CD_{5}\in(0.5,\penalty 10000\ 0.75]$; and Panel (c) uses $CD_{5}\in(0.75,\penalty 10000\ 1]$. Each panel compares the Park et al. (2023) methodology to our (III) methodology. The respective horizontal lines show the 1980 values for comparison purposes. The vertical line at 2010 represents the end of the Park et al. (2023) analysis.
Where do the differences come from? Fig. 6 illustrates that the Park et al. (2023) methodology results in declining CD indices over 1980-2010 in each highly disruptive patent group—particularly for $CD_{5}\in(0.75,\penalty 10000\ 1]$. [^14] In particular, the absolute number of patents in 2010 was 84% of the patents in 1980 in $CD_{5}\in(0.25,\penalty 10000\ 0.5]$, 71% in $CD_{5}\in(0.5,\penalty 10000\ 0.75]$, and 47% in $CD_{5}\in(0.75,\penalty 10000\ 1]$. A sensible conclusion based on their methodology is that the number of highly disruptive patents declined over 1980-2010.151515One potential reason why the authors suggest ”remarkable stability in the absolute number of highly disruptive works”(p. 140) may be that Figure 4b (p. 141) in Park et al. (2023) also shows $CD_{5}\in(0,\penalty 10000\ 0.25]$), which includes a substantially larger number of patents but does not belong in the highly disruptive category (according to their definition). As a result, the vertical axis masks the declines in the other three patent groups. Once truncation bias and exclusion bias are corrected for, however, Fig. 6 reveals that the number of highly disruptive patents increased over 1980-2010 in all three patent groups. [^16] In all three panels, the absolute number of highly disruptive patents differs from that reported in Fig. 4b (p. 141) of Park et al. (2023) because the authors only consider a subset of technologies (see Appendix A). Apart from this difference, the shape of the change is the same. Extending the data to 2016, moreover, indicates that the number of very highly disruptive patents more than doubled and the number of patents in the other disruptive patent groups increased significantly. This analysis also confirms Methodology (I) distortions—especially for very highly disruptive patents—as shown in the right panel of Fig. 6: the left-side of the U-shape arises from truncation bias (i.e., no citations to pre-1976 patent grants); the right-side arises from exclusion bias (i.e., no citations to patent applications).
@@
Appendix Table B.1 examines truncation bias in more detail. This table shows how patents in 1980 that belong to a particular $CD_{5}$ category using the Park et al. (2023) methodology (with truncation bias) are distributed across $CD_{5}$ categories using our methodology (without truncation bias). Backward citation truncation of pre-1976 patents misclassifies a relatively high proportion of patents as highly disruptive. For example, 25 percent of the USPTO patents in 1980 that are considered very highly disruptive with truncation bias are actually consolidating without truncation bias.
@@
Appendix Table B.2 examines exclusion bias in more detail. This table shows how patents in 2016 that belong to a particular $CD_{5}$ category using the Park et al. (2023) methodology (with exclusion bias) are distributed across $CD_{5}$ categories using our methodology (without exclusion bias). Backward citation exclusion to patent applications similarly misclassifies a relatively high percentageof recent patents as highly disruptive. For example, 19 percent of the USPTO patents in 2016 that are considered very highly disruptive using the Park et al. (2023) methodology are actually consolidating. Determining the disruptive potential of individual patents thus requires careful consideration to ensure against measurement bias and misclassification. [^17] Note that our findings also affect variations of the index proposed in Park et al. (2023) which are partly based upon other research Leydesdorff et al. (2021); Bornmann et al. (2020). In Appendix Fig. B.8 and B.9, we show the time series for several variations of the CD index used by Park et al. (2023): e.g., with respect to an extended time-window for forward citations or to a normalization of the index. In all cases, the truncation of backward citations leads to an analogous upward bias as described above.
@@ 7 Discussion
Our examination of the Park et al. (2023) CD index methodology was prompted by the stark decrease in the share of disruptive innovation in science and technology reported by the authors in patents over 1980-2010 and in scientific publications over 1945-2010. Our methodology and analysis of patents suggest results that either question or differ markedly from the Park et al. (2023) results, in at least four important respects.
First, we find that the secular decline in the average CD index found by Park et al. (2023) is mainly due to the truncation of backward citations, while a modest increase in the average CD index post-2005 is mainly due to the exclusion of backward citations to patent applications. In comparison to the Park et al. (2023) methodology, correcting for both biases results in a CD index that starts significantly lower in 1980, modestly declines to 2005, and is largely stable since. A CD index without measurement bias exhibits a marginal decline (viz., two and a half percentage points) but no secular decline (viz., 17 percentage points) in disruptive innovation over 1980-2010, as suggested by the authors. Second, we find that the starting truncation date does not matter. In particular, any backward citation cutoff in any chosen year produces similar CD index patterns as suggested by Park et al. (2023). Third, we find that the size of the truncation bias differs markedly across aggregate WIPO technology classes, given differences in the pattern and pace of innovation among these technologies. In particular, it is most pronounced in Mechanical and least pronounced in IT, precisely because more backward citations are truncated in Mechanical and less backward citations are truncated in IT. Fourth, we find that the number of highly disruptive patents has notably increased—particularly since 2008. Once truncation bias and exclusion bias are corrected for, our results challenge Park et al. (2023) in their assessment of stability in the number of highly disruptive patents over time.
@@
We suggest that our analysis is important for the research community that may work with biased data or overlook measurement issues—either via the CD index or other related measures. Our analysis highlights, in particular, the dangers of truncation bias from backward citations and exclusion bias from citations to patent applications. Our analysis suggests, perhaps more importantly, caution against overt policy interventions, regulatory changes, or substantive industry and firm reorganizations that seek to improve or disrupt the innovation status-quo based upon the Park et al. (2023) results. If it was accurate that breakthroughs in patented scientific knowledge have substantially decreased over time, then the current organization of the entire technology innovation process should be reevaluated. In particular, reforms that seek to increase disruptive innovation or restart innovation within R&D centers and research labs—e.g., in government agencies, universities and firms—would be deemed necessary and paramount. But changes to the current science and innovation process without a solid understanding of the facts in evidence is wrongheaded and might have unintended consequences.
@@ 9 Declaration of competing interest
10 Acknowledgments
We thank the editor, Adam B. Jaffe, and three anonymous referees for valuable comments that helped improve this paper.
@@ A Data
The study uses data from the February 21, 2023 version of PatentsView, which includes all U.S. patents published between 1976 and September 29, 2022. As in previous research Park et al. (2023); Funk and Owen-Smith (2017), we focus only on USPTO patents. We include PatentsView data that assigns each USPTO patent to technology categories as defined by the World Intellectual Property Organization (WIPO) Schmoch (2008). The specific WIPO technologies used to create our aggregated technology groups are provided in Appendix Table B.3. It is important to note that our technology classification differs from that of Park et al. (2023), who categorize patents based on NBER technology fields that are no longer available.
@@
In Sections 2-4, we use all granted utility patents over 1980-2010. Park et al. (2023) examine granted utility patents over this timeframe and in the NBER technology fields of "Chemical,Computers and Communications,Drugs and Medical,Electrical and Electronic,änd "Mechanical."Their data includes 3,046,672 patents and 29,777,375 backward citations (see replication data file patentsview_analytical_df.csvpatentsview\_analytical\_df.csv of Park et al. (2023)). We examine granted utility patents over the same timeframe, but cover all technology fields. Our data thus includes 3,662,051 patents and 42,617,016 backward citations in the case of no truncation (i.e., as they appear in the original patent documents) and 35,192,186 backward citations in the case of truncation.
@@
In Sections 5-6, we expand the dataset over 1980-2016, which results in 5,324,224 granted utility patents and 74,943,850 backward citations. With the passage of the Inventor Protection Act of 1999, patents beginning in November 2000 include backward citations to patent applications and to patent grants. Our methodology (III) includes all backward citations to patent grants and to patent applications that were eventually granted using supplemental PatentsView data. We replace all citations to patent applications with the corresponding patent grant information: i.e., the grant number and publication date. This replacement can result in some backward citations with publication dates later than the citing patent, but the backward citation publication date is not relevant to the $CD_{5}$ index methodology. What matters are publication dates of forward citations. We do not include patent applications published within five years after the publication year of a focal patent but with corresponding patent grants published later in the analysis. This approach results in 87,880,452 backward citations.
The PatentsView citation table contains patents that cite the same patent using different "publication dates", mainly due to typos in the patent reference lists. We apply the following correction: For each granted USPTO patent, we use the reference list and link each patent mentioned to its original patent document, using the patent number indicated. This provides the publication date for each cited patent as it appears in the original published patent document. We replace the date in the reference list with the date in the original document in all cases where the dates differ. We do the same for citations to patent applications. This correction is only possible for citations to patents published after 1975, as the original pre-1976 patent documents are not available in PatentsView. We therefore cannot correct the date of backward citations to patents published before 1976. We note that backward citation date is irrelevant for the CD index calculation, however, because it is not tied to a specific observation period. An incorrect date could nevertheless be problematic for forward citations where timing is important. Since we compute the CD index only for patents from 1976 onwards (and in most cases from 1980 onwards), only forward citations to patents published after 1975 are relevant. For all these citations, we correct the publication date provided by PatentsView using the procedure described above.